# UniSage: A Unified and Post-Analysis-Aware Sampling for Microservices

### v1

@@ 5.5. RQ2: Effectiveness of root cause analysis
UniSage 0.411 0.732 0.589 0.545 0.606 0.620
Table 4 reports the downstream performance of various RCA methods (i.e., MicroRank, TraceRCA, TraceAnomaly, and PDiagnosis) under different sampling strategies on two datasets. For fairness, we compare existing approaches strictly within the conventional sample-before-analysis paradigm, where the RCA models operate only on the sampled data. In contrast, our analysis-before-sample approach leverages RCA results derived from the full data stream to guide sampling. This design not only demonstrates the effectiveness of our method but also highlights the superiority of the proposed post-analysis-aware paradigm.
Sampling quality strongly affects RCA effectiveness. From the results under the sample-before-analysis paradigm in Table 4, we observe that higher-quality sampling strategies consistently lead to better RCA accuracy across methods and datasets. The Uniform baseline consistently performs the worst, with $AC@1$ close to zero. Sifter and Sieve show moderate improvements, but their performance remains limited due to their general-purpose design. TraStrainer, which leverages system metrics, performs better and ranks highest among baselines (e.g., $MRR$ of 0.421 on OnlineBoutique with MicroRank), yet it is still lower than UniSage (e.g., $MRR$ of 0.497 on OnlineBoutique with TraceRCA). These results emphasize that sampling strategy quality is critical for effective RCA, especially when the trace budget is constrained. Poorly chosen samples can omit anomaly-relevant data, undermining downstream analysis.
Post-analysis sampling yields superior RCA performance. From the analysis-before-sample paradigm results in Table 4, we observe that UniSage consistently achieves the highest accuracy across all RCA methods and datasets. Unlike baselines that sample first and analyze later, our approach first conducts RCA on the full dataset and then guides sampling with the results, ensuring high coverage of anomaly-relevant data. For instance, with MicroRank, UniSage attains $AC@1$ scores of 0.411 and 0.545 on OnlineBoutique and TrainTicket, respectively, substantially outperforming the best baseline TraStrainer ($AC@1$ of 0.339 and 0.333).
@@
5.6. RQ3: Efficiency and Scalability
@@
(a) Comparison of sampling efficiency between UniSage and trace-only baselines with Uniform log sampling.
(b) Proportional time consumption of UniSage.
Figure 7. Efficiency and proportional time consumption results.
We assess the runtime efficiency and scalability of different sampling strategies by measuring their total sampling time on both datasets (Table 2). For a fair comparison, we add the time of Uniform log sampling (i.e., random selection) to all trace-only baselines (i.e., Sieve, Sifter, TraStrainer, Uniform), as UniSage uniquely supports joint sampling of both traces and logs. Note that UniSage ’s total time also includes RCA, as it follows a post-analysis sampling strategy, unlike baselines, which measure sampling time only.
As shown in Figure 7(a), UniSage achieves a strong balance between efficiency and scalability. On the large-scale OnlineBoutique dataset (nearly 40x more traces and 19x more logs), UniSage completes end-to-end analysis in 739.4 seconds, significantly faster than Sieve (3,364.1s) and TraStrainer (1,160s), while offering much stronger performance. On the smaller TrainTicket dataset, UniSage finishes in 142.3 seconds (4.3s per case), remaining competitive despite the added RCA overhead. In contrast, some baselines, such as Sifter, show fast runtime on smaller datasets (18.6s), but do not scale well to larger workloads. These results demonstrate that UniSage not only integrates trace and log sampling effectively, but also remains practical for real-world deployment at scale.
Time Consumption Proportion. To better understand UniSage’s internal overhead, we divide its runtime into three stages: log sampling, trace sampling, and RCA. As shown in Figure 7(b), RCA accounts for 49% of the total time, trace sampling for 47%, and log sampling for only 4%. This breakdown demonstrates that our RCA module is relatively lightweight, despite being executed before sampling. It effectively addresses a key concern of the analysis-before-sample paradigm, the risk of high analysis latency hindering practical deployment. In contrast, UniSage maintains a balanced runtime profile suitable for real-world usage.
@@ 6.1. Limitations
Most recent research on trace sampling, notably HindSight (hindsight, ), introduces the concept of retroactive sampling, where trace data is lazily retrieved to perform biased sampling earlier in the trace lifecycle. This approach improves the efficiency of biased sampling by deferring decisions until more information is available. While HindSight shares a similar motivation with UniSage (i.e. performing analysis before sampling), it focuses solely on edge-case traces and does not support downstream diagnosis tasks. In contrast, UniSage not only considers edge-cases but also emphasizes anomaly relevance and diagnosis support, enabling the retention of more informative and actionable data. Another limitation lies in the reliance on lightweight anomaly detection and RCA algorithms in the initial analysis stage. While these components are efficient and scalable, their accuracy may impact the quality of sampling decisions. Future work could investigate how more advanced or adaptive analysis techniques affect the sampling outcome.
6.2. Generalizability
While our current evaluation focuses on AD and RCA, we believe UniSage can generalize well to other downstream tasks such as performance profiling and workload characterization. For performance profiling, UniSage’s sampling strategy incorporates performance indicators (e.g., trace durations, metrics), which are commonly used in profiling tasks. This enables the retained data to still reflect key performance behaviors, even under storage constraints. For workload characterization, which aims to comprehensively identify workload patterns, UniSage is explicitly designed to capture diverse and rare behaviors (e.g., anomalies, edge cases). This makes it particularly suitable for characterizing workloads when observability data is limited. While our current evaluation focuses on AD and RCA, we agree that extending empirical validation to other tasks is valuable and plan to explore this in future work.
@@ 6.3. Threats to Validity
The main threat to internal validity lies in the implementation of baselines. For Sieve and TraStrainer, we use their publicly available open-source implementations. However, Sifter lacks open-source code, so we carefully re-implement it based on descriptions in the original papers, ensuring the use of identical libraries and settings whenever possible.
The primary threat to external validity concerns the generalizability of our results. Although we evaluate UniSage on two widely-used microservice benchmarks, the number of injected fault types remains limited (five in total), which may not capture the full diversity of real-world failure modes. Additionally, while some production datasets exist, they often include only traces without corresponding logs or metrics, preventing a comprehensive evaluation of our unified sampling strategy. To further validate the robustness of UniSage, future work could explore its application in more diverse system architectures, real production environments, and under more complex fault scenarios.
7. Related Works
Distributed tracing is a fundamental technique for monitoring and diagnosing the behavior of distributed systems. It captures the end-to-end execution of requests across services, providing visibility into performance bottlenecks and failure points. This is essential for identifying issues that are nondeterministic or difficult to reproduce in isolation. By tracing the flow of a request through different components, developers can better understand and debug complex distributed systems. Over the years, significant progress has been made in tracing frameworks. X-Trace (X-Trace, ) focuses on tracing across multiple system layers, enabling detailed visibility into application behavior. Dapper (dapper, ), Google’s tracing system, demonstrates the viability of large-scale distributed tracing in production environments. Canopy (canopy, ), developed at Facebook, captures performance data and causal relationships across entire request paths to support end-to-end analysis. In recent years, many open-source tracing frameworks such as Jaeger (Jaeger, ) and Zipkin (Zipkin, ) have been widely adopted in practice. Distributed tracing has enabled a wide range of system observability and analysis tasks, including performance profiling (mi2013toward, ; sambasivan2011diagnosing, ), anomaly detection (microhecl, ; tracecrl, ; nedelkoski2019anomaly, ), and failure diagnosis (gan2021sage, ; Eadro, ; Nezha, ; MRCA, ; li2025coca, ; liu2020unsupervised, ; hemirca, ; wu2025trace, ). As trace volume grows rapidly in production environments, several trace reduction techniques have been proposed (tracemesh, ; steam, ; huang2021sieve, ; sifter, ), with head sampling (canopy, ; dapper, ) and tail sampling (huang2021sieve, ; sifter, ; las2018weighted, ) being the most commonly adopted strategies.
Recent studies have started to incorporate troubleshooting and anomaly-awareness into sampling decisions. For example, TraStrainer (TraStrainer, ) introduces a trace sampling approach that adapts dynamically to system runtime state by leveraging metrics. By jointly considering trace diversity and metric anomalies, it prioritizes performance-relevant traces. However, TraStrainer’s detection scope is limited to anomalies that manifest in metric signals, leaving out other types of failures, and it only reduces traces. Hindsight (hindsight, ) presents a retroactive sampling framework targeting edge-case request diagnostics. Although it enhances sampling precision by using full-trace hindsight, it remains focused solely on traces and request-level analysis. In contrast, our work proposes a unified post-detection sampling framework that supports both traces and logs. By incorporating signals from metrics, logs, and traces, our system can detect and retain a broader range of anomaly types. Furthermore, our method incorporate interpretable signals that explain the rationale behind each sampling decision, revealing why the retained data is anomaly-related or represents uncommon behavior. This design effectively bridges the gap between fault diagnosis and data reduction.
@@ 8. Conclusion
In this paper, we presented UniSage, a novel post-analysis-aware sampling framework for observability data in distributed systems. Unlike traditional approaches that sample data prior to analysis, UniSage reverses this process by first performing lightweight anomaly detection and root cause analysis, and then retaining a concise yet informative subset of data for future use. Our method jointly considers both traces and logs, and emphasizes anomaly relevance and data diversity, ensuring that the preserved samples are not only valuable for fault diagnosis but also interpretable. Moreover, by addressing the previously overlooked problem of log sampling and offering a unified solution, UniSage fills an important gap in the observability pipeline. Extensive experiments on two widely-used microservice benchmarks demonstrate that UniSage outperforms existing sampling strategies in terms of effectiveness, efficiency, and support for root cause analysis. In future work, we plan to explore adaptive sampling strategies that dynamically adjust retention policies based on evolving system behavior, as well as extend our framework to support other modalities such as events and alerts.

### v2

@@ 4.5. RQ2: Effectiveness of Paradigm Shift
Boutique UniSage 0.411 0.732 0.589 0.372 0.713 0.843 0.626 0.807 0.943
UniSage $w/o~\mathcal{A}$ 0.411 0.732 0.589 0.205 0.422 0.547 0.347 0.539 0.774
[…]
AIOps21 UniSage 0.315 0.528 0.474 0.897 1 1 - - -
UniSage $w/o~\mathcal{A}$ 0.315 0.528 0.474 0.420 0.562 0.650 - - -
[…]
4.6. RQ3: Ablation Study
Table 5 validates the design of our dual-pillar sampling and multi-modal analysis by comparing UniSage against four variants.
[…]
Answer to RQ3: The results confirm that multi-modal integration is essential for accurate RCA and sampling, while the dual-pillar strategy is indispensable for ensuring comprehensive sampling coverage.
4.7. RQ4: Efficiency and Scalability
@@
(a) Time cost of UniSage and trace sampling baselines with uniform log sampling.
(b) proportional time consumption of different components in UniSage.
Figure 8. Efficiency and time breakdown results.
We assess runtime efficiency on three datasets: TrainTicket (small), Boutique (large), and AIOps21 (real-world). To ensure a fair comparison with UniSage’s joint sampling, we add the time of uniform log sampling to all trace-only baselines (Sieve, Sifter, TraStrainer, Uniform). Notably, UniSage’s reported time includes the complete post-analysis RCA process, whereas baselines measure only sampling overhead.
Time cost. As shown in Figure 8(a), UniSage demonstrates superior scalability. On the large-scale Boutique dataset, it requires only 13.2s per case, substantially outperforming Sieve (60.1s) and TraStrainer (20.7s). Similarly, on the high-volume AIOps21 dataset (5.3M logs), UniSage finishes in 10.1s, approximately $4.4\times$ and $2.5\times$ faster than Sieve and TraStrainer, respectively. While simple baselines like Sifter are faster on the small TrainTicket workload (0.6s), they fail to scale, with latency increasing $13\times$ on larger datasets. In contrast, UniSage maintains stable performance as data complexity grows. Time breakdown. Figure 8(b) decomposes UniSage’s runtime into RCA (49%), trace sampling (47%), and log sampling (4%). This confirms that the RCA module is lightweight, effectively addressing concerns that post-analysis induces prohibitive latency and validating UniSage’s suitability for real-world deployment.
Answer to RQ4: UniSage achieves strong sampling capability with minimal analysis overhead, remaining efficient and scalable even when including post-analysis. This confirms the production viability of analysis-aware sampling.
5. Related Works
Distributed tracing captures the end-to-end execution of requests, providing essential visibility into performance bottlenecks and nondeterministic failures in distributed systems. Foundational frameworks include X-Trace (Fonseca et al., 2007) for multi-layer visibility and Dapper (Sigelman et al., 2010), which demonstrated the viability of large-scale tracing. Facebook’s Canopy (Kaldor et al., 2017) further advanced end-to-end analysis by capturing causal relationships. Recently, open-source frameworks like Jaeger (Jaeger, 2025) and Zipkin (Zipkin, 2025) have seen widespread adoption. Tracing has enabled a wide range of system observability and analysis tasks, including performance profiling (Mi et al., 2013; Sambasivan et al., 2011), anomaly detection (Liu et al., 2021; Zhang et al., 2022; Nedelkoski et al., 2019), and failure diagnosis (Gan et al., 2021; Lee et al., 2023a; Yu et al., 2023b; Wang et al., 2024; Liu et al., 2020a; Zhu et al., 2024). To manage rapidly growing data volumes, several trace reduction techniques have been proposed (Chen et al., 2024; He et al., 2023; Huang et al., 2021; Las-Casas et al., 2019), with head sampling (Kaldor et al., 2017; Sigelman et al., 2010) and tail sampling (Huang et al., 2021; Las-Casas et al., 2019, 2018) being the most commonly adopted strategies.
Recent studies have begun incorporating anomaly awareness into sampling. TraStrainer (Huang et al., 2024a) dynamically prioritizes traces based on metric anomalies but misses failures undetectable via metrics and is limited to trace data. Similarly, Hindsight (Zhang et al., 2023) offers retroactive sampling for edge-case diagnostics but remains restricted to request-level trace analysis. In contrast, we propose a unified, post-analysis-aware sampling framework for both traces and logs. By integrating signals from metrics, logs, and traces, our approach captures a broader spectrum of anomalies. Moreover, unlike black-box methods, UniSage provides explicit statistical indicators to explain sampling decisions, effectively bridging the gap between fault diagnosis and data reduction.
@@ 6.1. Limitations
UniSage adopts a tail sampling strategy, which is inherently biased toward informative samples and therefore incurs higher computational overhead than head sampling methods based on random selection. Nevertheless, our overhead remains fully acceptable, demonstrating the practicality of UniSage. An interesting direction for future work is to explore hybrid sampling strategies that integrate head and tail sampling, enabling a principled trade-off between efficiency and effectiveness. Moreover, the primary focus of this work is to examine whether the proposed paradigm shift leads to better sampling decisions, rather than to incorporate advanced diagnostic techniques requiring training or modeling. Accordingly, we employ simple yet effective detection methods, such as the $k-\sigma$ rule and PCA. Future work could investigate how more sophisticated or adaptive analysis techniques influence the quality and robustness of the resulting sampling decisions.
@@ 6.2. Threats to Validity
The main threat to internal validity lies in the implementation of baselines. For Sieve and TraStrainer, we use their publicly available open-source implementations. However, Sifter lacks open-source code, so we carefully re-implement it based on descriptions in the original paper, ensuring the use of identical libraries and settings whenever possible. Threats to external validity mainly stem from the representativeness of the evaluated datasets. Although our evaluation includes one real-world industrial dataset and two widely used open-source microservice benchmarks, the diversity of fault types remains limited (7 in total). Real-world microservice systems may exhibit more complex and diverse failure patterns, which could affect the generalizability of our findings. Additionally, the industrial dataset (AIOps21) contains only 14 microservices, whereas production systems in large-scale enterprises may involve more services; large-scale industrial datasets are typically unavailable due to confidentiality constraints. However, our efficiency evaluation demonstrates that the runtime of UniSage scales linearly as the dataset size and system complexity increase, partially mitigating this concern.
@@ 7. Conclusion
In this paper, we addressed the critical challenge of balancing data volume with diagnostic utility in microservice observability. We introduced UniSage, a unified sampling framework that fundamentally shifts the paradigm from sample-before-analysis to a post-analysis-aware strategy. By integrating lightweight anomaly detection and root cause analysis prior to sampling, UniSage effectively breaks the trade-off between storage overhead and the retention of critical diagnostic context. Our dual-pillar strategy, which combining analysis-guided sampling for anomalies and edge-case sampling for diversity, demonstrated superior performance in our evaluations. At a sampling rate of 2.5%, UniSage successfully retained 71% of critical traces and 96.25% of relevant logs, significantly outperforming state-of-the-art methods which often discard vital diagnostic data. Furthermore, with an average processing time of just 10 seconds for 20-minute data blocks, UniSage proves to be a highly efficient and practical solution for real-world production environments. We believe that UniSage paves the way for more intelligent, context-aware observability systems that maximize insight without imposing prohibitive computational burdens.